Imagine you are planning to use circles to visually compare the populations of California (38,800,000), Texas (28,240,245), and New York (19,807,500). The size of each population will be represented by the area of each circle, not the radius. Round your answer to the nearest tenth.
If the circle for New York will have a radius of 5 cm, what would be the size (area) of that circle (which represents the population of New York)?
____ cm^2

Found 3 solutions by t0hierry, ikleyn, Theo:
Answer by t0hierry(194)

r = 5
A = pi r^2
A = pi * 25 = 75.5 cm^2

Answer by ikleyn(52847)

.
r = 5

A = pi * r^2

A = pi * 25 = 3.14159*25 = 78.53975 cm^2 (approximately), or 78.5 cm^2 rounded to the nearest tenth.


------------

Regarding the answer by @t0hierry, I assume that a physicist with a PhD degree could make his calculations in a more accurate manner . . .



Answer by Theo(13342)

area of a circle is pi * r^2

if the radius of the circle for new york is 5 centimeters, then the area of the circle for new york is pi * 5^2 = 25 * pi.

if you let N equal the population for new york and T equal the population for texas and C equal the population for california, then:

the area of the circle for texas would be T / N * 25 * pi.

the area of the circle for california would be C / N * 25 * pi.

if you want to derive the radius of the circle, the formula for that is radius = square root of (area / pi).

for example, the radius for the area of the circle that represents new york is equal to the square root of (25 * pi / pi) = square root of (25) = 5.

i did the calculations in excel.

these are the results.

$$$

i'm not exactly sure what you are looking for, but if it's the area of the circle that represents new york, then your answer is 25 * pi square centimeters which is equal to 78.5 square centimeters rounded to the nearest tenth.